The action level for lead (Pb) in water is 0.015 ppm. What is the concentration in molarity (M, mol/L)? Respond with the correct number of significant figures in scientific notation.
Answer #1

ppm represent concentration of Pb in mg/L . Thus,

0.015 ppm = 0.015 mg/L = 0.015×10-3 g/L

Molar mass of Pb = 207.2 g/mol

Molarity = (0.015×10-3)/207.2 mol/L = 0.72×10-7 mol/L

Please give feedback!